March 14, 2011

Observers: J. Bechtold, D. Thompson, F. Bian, L. Jiang

2:25pm. Lights off in the dome. Started series of 900s darks for Linhua's observations tonight.

5:15pm. Taking internal flats with Halo 3 for imaging

5 exposures in J, 6 sec each starting with file 5

5 exposures in Ks, 5 sec each

5 exposures in H, 4 sec each

frames 21-40 Dave did WCS testing

5:51pm. Taking sky flats (thin clouds)

10 exposures in Br_gam at 5:54 pm, exposure time 2s, counts ~11000, images 41-50

10 exposures in Br_gam at 6:00 pm, exposure time 2s, counts ~7000, images 51-60

10 exposures in H2 at 6:01 pm, exposure time 2s, counts ~8500, images 61-70

counts are getting higher due to clouds, stopped for a few minutes

Alt changed to 60 deg

Cirrus everywhere as sun is setting. Twilight flats are problematic.

Checked telescope center on Lucifer detector -- ok

6:45pm. Going to star near 3C220 to collimate. Still light out, clouds moving in.

7:10pm. At 3C220 Too bright out to see guide star.

7:20pm. Not collimating.

7:25pm. Collimated

Executing 3c220.acq.txt

7:45pm Starting integration. 3c220.spec.txt starting with file #79 seeing 1.8-2.2"

LOTS OF CIRRUS

8pm. Seeing 1.1" or so, but still cloudy

8:05pm. Lots of clouds, but seeing 0.8" in guider. However we're getting yellow warnings from the WFS and guide star is elongated. No wind.

9:05pm. Started 3c220.spec2.txt, file starts 94. Seeing 1" still cloudy. This script is the same as spec1 except it starts with an offset to get back to the center.

9:50pm. Seeing 2.2" Looks like it might get clearer eventually.

10:20pm Seeing 1.5"

10:30pm Doing 3c220.arcs.txt

10:35pm Doing 3c220.flat.txt

10:35pm 3c220_tell2.acq.txt Going to the telluric standard

10:50pm 3c220_tell2.spec.txt Seeing 1" Saturated

10:55pm qq.txt Changed exposure time to Dit=5, NDIT=4

11:00 pm Executing sdss1343.slit2.acq.txt

11:20pm Executing sdss1343.spec.txt seeing 1.2"

aborted after one image because AGW unit had RPC communication error and we stopped guiding

11:35pm Reaquiring with sdss1343.slit2.acq.txt. seeing .8" in guider.

11:45pm Execute sdss1343.spec.txt seeing 1.0"

1:05am sdss1343.arcs.txt

1:10am sdss1343.flat.txt

1:15am sdss1343_tell2_slit2.acq.txt

1:20am sdss1343_tell_spec.txt

1:25am sdss1343_tell2_slit1.acq.txt

1:30am sdss1343_tell_spec.txt seeing .5"

1:35am sdss1343_slit1.acq.txt Tonight the central wavelength will be at 1.88microns

1:45am sdss1343.spec.txt seeing in guider .45"

3:10am sdss1343.arcs.txt for slit 1

3:15am sdss1343.flat.txt for slit 1

3:16am executing jiang_acq.sdf66.txt Still hazy.

3:35am executing jiang_sci_acq.txt Seeing 0.58" in guider

5:45am. aborted jiang_sci_acq.txt and executed jiang_arc_sdf66.txt and jian_flat_sdf66.txt
Topic revision: r6 - 14 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback